OpenStudy (anonymous):

Complete the following ordered pairs for the equation? 9x + -5y = 60 (-5, _) (_, -3) (10, _)

OpenStudy (callisto):

For (-5, _), put x=-5 into the equation. 9(-5) + -5y = 60 Now, solve y and put the value in that blank. For (_, -3) , put y=-3 into the equation. 9x + -5(-3) = 60 Now, solve x and put the value in that blank. For (10, _), it's similar to (-5, _).

OpenStudy (callisto):

Can you do it?

OpenStudy (anonymous):

I have no idea how to do any of this, so no. lol

OpenStudy (callisto):

For the first one, 9(-5) + -5y = 60 First, what is 9(-5) ?

OpenStudy (anonymous):

-45

OpenStudy (callisto):

So, you get -45 - 5y = 60 Got it so far?

OpenStudy (anonymous):

I think so.

OpenStudy (callisto):

Now, add 45 to both sides, what do you get?

OpenStudy (anonymous):

Both sides of what...? lol. I'm sorry teaching me math isn't simple. I really don't understand any of it.

OpenStudy (callisto):

Both sides of -45 - 5y = 60

OpenStudy (anonymous):

I have no idea what you're talking about... :(

OpenStudy (callisto):

Add 45 to both sides of -45 - 5y = 60: -45 - 5y +45= 60 + 45 Now, simplify it, what do you get?

OpenStudy (anonymous):

-5 and -21? I DON'T KNOW!

OpenStudy (callisto):

What is 60 + 45?

OpenStudy (anonymous):

105

OpenStudy (callisto):

What is -45 - 5y +45?

OpenStudy (anonymous):

180?

OpenStudy (callisto):

Not really.. -45 - 5y +45 = -45 + 45- 5y = ...?

OpenStudy (anonymous):

Ugh, I have no idea... I'm so lost...

OpenStudy (callisto):

That is -45+45?

OpenStudy (anonymous):

0

OpenStudy (callisto):

-45 - 5y +45 = -45 + 45- 5y = 0 - 5y = -5y Agree?

OpenStudy (anonymous):

I guess, I don't know what just happened.. I'm sorry.

OpenStudy (callisto):

Do you know what happened now?

OpenStudy (anonymous):

Honestly, no. I don't.

OpenStudy (callisto):

Very honest :) We need to solve the equation :9(-5) + -5y = 60 Since the left side and right side are equal, you can find the value of y. For example, if you have a equation x+1 = 4, and you need to find x, just think what number plus one is equal to four. And that number is x. Got this example?

OpenStudy (anonymous):

Yeah I understand that one. It would be 3 right?

OpenStudy (callisto):

Yes! Actually, don't be frightened when you see that equation 9(-5) + -5y = 60 ! When you simplify and apply that concept, you can solve it easily too! Now, take a look at the equation 9(-5) + -5y = 60 As you've mentioned, 9(-5) = -45 So, the first term on the left is equal to -45. Rewrite the equation and you'll get: -45 + (-5y) = 60 Got it?

OpenStudy (anonymous):

See no, that's where I get lost..

OpenStudy (callisto):

Which part??

OpenStudy (anonymous):

All of it. Thanks for your help, but I'm just gonna give up..

OpenStudy (callisto):

9(-5) = -45 Agree?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

 Got it?